OpenStudy (anonymous):

If a particular radio telescope is 100ft in diameter and has a cross section modeled by the equation x^2=167y, how deep is the parabolic dish? What is the location of the focus? can someone help me with this?

OpenStudy (amistre64):

learned a bit more about this stuff today i did....

OpenStudy (amistre64):

x^2 = 4ay where a is the distance from the vertex..... and your vertex is at 0,0

OpenStudy (amistre64):

167/4 is a so the focus is at..... 0+ 167/4

OpenStudy (anonymous):

right. i have that its the first part that im having trouble with

OpenStudy (amistre64):

ok.....we know that it 100 feet across, so that is 50 feet in both directions, when y = 50, what is x^2 = ?

OpenStudy (amistre64):

is it headed in the x direction or the y?

OpenStudy (anonymous):

the parabola opens up

OpenStudy (amistre64):

then I would turn it into the normal way of viewing this: y =(1/167)x^2. when x = 50, whats y=?

OpenStudy (amistre64):

2500/167...

OpenStudy (anonymous):

so thats how deep it is?

OpenStudy (amistre64):

yup,...

OpenStudy (amistre64):

the numbers are the same, we just repositioned them into the "normal" way that this equation would look.

OpenStudy (amistre64):

the vertex is at (0,0) thats half way, half of the diameter is 50...plug it in :)

OpenStudy (amistre64):

14.97 is what I get

OpenStudy (anonymous):

k I'll try that. thanks!

OpenStudy (amistre64):

youre welcome :)

OpenStudy (amistre64):

have you ever worked with quadratic equations?

OpenStudy (amistre64):

cause thats all a parabola is...

OpenStudy (anonymous):

yes. im in pre calc at a jr college, this stuff is being thrown at me quick but im doing my best.

OpenStudy (amistre64):

phcc sound familiar to you?

OpenStudy (anonymous):

no whats that?

OpenStudy (amistre64):

its the community college I go to right now :) just wondering if you were in my trig/pre calc class ....

OpenStudy (anonymous):

na im at tcc

OpenStudy (amistre64):

well, good luck to ya :)

OpenStudy (anonymous):

thanks!